# GraphPerf-RT: A Graph-Driven Performance Model for Hardware-Aware Scheduling of OpenMP Codes

### v2

@@ 4 Design Methodology
This section presents the GraphPerf-RT architecture following the problem formulation established in Section 3.1. The model takes as input the task DAG $G=(V,E)$, device sheet $D$, runtime state $\mathcal{S}$, and scheduling action $a=(m,f,p)$, and outputs performance predictions $y$ with calibrated uncertainty $\mathcal{U}$ for each target metric.
GraphPerf-RT Design Pipeline Inputs: application graph $G=(V,E)$; hardware config $\mathcal{H}$; system state $\mathcal{S}$ Nodes: Task (CFG feat., static/dyn. feats), Resource (core type, $f$, $T$), Memory (levels, BW/lat.) Type-specific encoders: MLPT, MLPR, MLPM Heterogeneous GAT (3–6 layers); edge types: $E_{TT},E_{TR},E_{RR},E_{RM}$; topo encodings Hierarchical pooling: POOL($V_{T}$) $\parallel$ POOL($V_{R}$) $\parallel$ POOL($V_{M}$) $\rightarrow$ graph embedding $\mathbf{h}_{G}$ Multi-task evidential heads $\rightarrow$ metrics $\{\text{time},\text{energy},\text{cache},\text{branch},\text{util}\}$ with NIG params $\{\gamma,\nu,\alpha,\beta\}$ Training: (1) feature learning $\rightarrow$ (2) evidential fine-tuning $\rightarrow$ (3) calibration
Figure 1: End-to-end pipeline.
@@ 4.1 Heterogeneous Graph Representation
Intuition: Why Task-Resource-Memory Decomposition?
Performance of parallel applications on heterogeneous SoCs arises from three interacting factors: (1) what computation runs (task structure, dependencies, code complexity), (2) where it runs (core types, frequencies, thermal state), and (3) how data flows (cache hierarchies, memory bandwidth, contention). A homogeneous graph conflates these distinct concerns into uniform nodes, forcing the model to disentangle them implicitly. Our heterogeneous decomposition makes these factors explicit: Task nodes capture application semantics (CFG complexity, DAG structure, data dependencies), Resource nodes capture hardware state (DVFS settings, utilization, thermals), and Memory nodes capture the memory hierarchy (cache levels, bandwidth constraints). The typed edges then encode the interactions that determine performance: task-task edges model parallelism and synchronization overhead, task-resource edges model scheduling decisions and affinity, resource-resource edges model shared cache contention between cores, and resource-memory edges model bandwidth bottlenecks. This explicit structure allows the GNN to learn interpretable attention patterns, for example, attending strongly to thermal headroom when predicting energy, or to critical-path edges when predicting makespan, rather than discovering these relationships from scratch in a homogeneous representation.
GraphPerf-RT uses a unified heterogeneous graph representation that captures both the application structure and the hardware architecture in a manner consistent with the abstraction in Section 3.2. This representation consists of three types of nodes ($V_{T}$, $V_{R}$, $V_{M}$) and multiple edge types ($E_{TT}$, $E_{TR}$, $E_{RR}$, $E_{RM}$) that encode different relationships within the system, allowing the model to reason about dependencies, placements, contentions, and resource constraints holistically.
4.1.1 Node Types
Table 1: Node types and their performance rationale.
Node Type Captures Why Separate?
$V_{T}$ (Task) What runs Code complexity, DAG structure, data deps
$V_{R}$ (Resource) Where it runs Core type, DVFS, thermal, utilization
$V_{M}$ (Memory) Data flow path Cache levels, bandwidth, contention
@@ 4.1.1 Node Types
Device constants from $D$ (e.g., DVFS tables, cache sizes) are broadcast as shared features across relevant nodes during encoding.
4.1.2 Edge Types
Table 2: Edge types and their performance rationale.
Edge Type Models Performance Impact
$E_{TT}$ (Task-Task) Parallelism, sync Critical path, fork/join overhead
$E_{TR}$ (Task-Resource) Scheduling Affinity, migration cost, load balance
$E_{RR}$ (Resource-Resource) HW topology Cache contention, cluster effects
$E_{RM}$ (Resource-Memory) Memory hierarchy Bandwidth bottlenecks, latency
This rich representation ensures the model captures cross-layer interactions (aligning with the data pipeline in Section 3.3) and enables the GNN to learn distinct attention patterns for each interaction type—for example, attending to $E_{TT}$ critical edges when predicting makespan, or to $E_{RR}$ sharing edges when predicting cache misses.
4.2 Graph Neural Network Architecture
GraphPerf-RT employs a GAT architecture specifically designed to handle heterogeneous graphs with multiple node and edge types, building on the encoders and backbone described in Section 3.4. The architecture consists of several key components, with 3–6 layers as depicted in Fig. 1, using hidden dimensions typically in $[128,256]$ for embeddings, and multi-head attention (e.g., 4–8 heads) to stabilize learning.
4.2.1 Type-Specific Encoders
@@
Since different node types have distinct feature spaces, we employ type-specific Multi-Layer Perceptrons (MLPs) to encode raw features into a common embedding space of dimension $d$ (e.g., 128):
4.2.2 Heterogeneous GAT Layers
The core of the architecture consists of multiple heterogeneous graph attention layers that aggregate information from neighboring nodes while accounting for edge types and their attributes. For each layer $l$, the update for node $v$ is:
where $\mathcal{R}=\{TT,TR,RR,RM\}$ is the set of edge types, $\mathcal{N}_{r}(v)$ is the neighborhood of $v$ under edge type $r$, $\alpha_{uv,r}^{(l)}$ is the attention coefficient, $\mathbf{W}_{r}^{(l)}$ is a type-specific linear transformation matrix, and $\sigma$ is a non-linearity (e.g., ELU [8]). Residual connections are added for deeper models to mitigate vanishing gradients.
4.2.3 Graph-Level Pooling
After the final graph attention layer $L$, we apply a hierarchical pooling strategy to obtain a graph-level representation. This involves separate pooling operations for different node types (e.g., mean, max, or attention-based pooling), followed by concatenation:
$\mathbf{h}_{G}=\text{CONCAT}(\text{POOL}_{T}(\{\mathbf{h}_{v}^{(L)}:v\in V_{T}\}),$
$\text{POOL}_{R}(\{\mathbf{h}_{v}^{(L)}:v\in V_{R}\}),\text{POOL}_{M}(\{\mathbf{h}_{v}^{(L)}:v\in V_{M}\}))$ (5)
@@ 4.3 Evidential Learning for Uncertainty Quantification
To provide uncertainty-aware predictions, GraphPerf-RT employs an evidential learning framework consistent with Section 3.4. Instead of directly predicting performance values, the model learns the parameters of a NIG distribution for each performance metric via multi-task heads on $\mathbf{h}_{G}$:
for each metric $k\in$ { makespan, energy, cache misses, branch misses, utilization }, where $\text{MLP}_{k}$ is a shared trunk (e.g., 2 layers, 128 units) with task-specific outputs (e.g., linear layers with softplus activations on $\nu_{k},\alpha_{k}>1,\beta_{k}>0$ to ensure valid distributions). The predictive distribution for metric $k$ is given by:
$p(y_{k}|\mathbf{h}_{G})=\text{NIG}(y_{k};\gamma_{k},\nu_{k},\alpha_{k},\beta_{k})=$
$\int\mathcal{N}(y_{k};\mu,\sigma^{2})\text{IG}(\sigma^{2};\alpha_{k},\beta_{k})\,d\sigma^{2}$ (7)
with $\mu\sim\mathcal{N}(\gamma_{k},\sigma^{2}/\nu_{k})$. The mean prediction and uncertainty estimates are derived as:
$\hat{y}_{k}$ $=\gamma_{k}$ (8)
$\text{Aleatoric Uncertainty}_{k}$ $=\frac{\beta_{k}}{\alpha_{k}-1}$ (9)
$\text{Total Variance}_{k}$ $=\text{Aleatoric}_{k}+\text{Epistemic}_{k}$ (11)
The model is trained by minimizing the negative log marginal likelihood of the evidential distribution, augmented with evidence regularization (to prevent underconfidence) and a ranking loss (e.g., pairwise on makespan to align with scheduling objectives):
$\mathcal{L}=\sum_{k=1}^{K}\sum_{i=1}^{N}[\frac{1}{2}\log(\frac{\pi}{\nu_{k,i}})$
$-\frac{\alpha_{k,i}-1}{2}\log(2\beta_{k,i}(1+\nu_{k,i}\epsilon_{k,i}^{2}))+$
$(\alpha_{k,i}+\frac{1}{2})\log(1+\nu_{k,i}\epsilon_{k,i}^{2})+$
$\log\frac{\Gamma(\alpha_{k,i})}{\Gamma(\alpha_{k,i}+1/2)}]+$
$\lambda\sum_{k}|\alpha_{k,i}-1|+\rho\mathcal{L}_{\text{rank}}$ (12)
where $\epsilon_{k,i}=(y_{k,i}-\gamma_{k,i})^{2}/(2\beta_{k,i})$, $\Gamma$ is the gamma function, $\lambda$ (e.g., 0.01) regularizes evidence strength, $\rho$ (e.g., 0.1) weights the ranking loss, and $\mathcal{L}_{\text{rank}}$ enforces ordering consistency (e.g., via margin-based pairwise loss on makespan predictions).
Handling Dynamic and Irregular Workloads.
@@ 4.4 Training Procedure
GraphPerf-RT is trained using a multi-stage procedure designed to handle the challenges of multi-task learning with uncertainty quantification, building on the data preprocessing in Section 3.3 (e.g., z-scoring per device, outlier filtering via MAD on residuals, 60/20/20 splits stratified by benchmark-input-mask-DVFS to avoid leakage):
Stage 1: Feature Learning. The model is first trained to learn effective representations using a standard multi-task regression loss (e.g., MSE per target, weighted by inverse variance), focusing on prediction accuracy without uncertainty quantification. Optimizer: AdamW with learning rate $10^{-3}$, batch size 32–128 (depending on graph size), epochs 50–100, with early stopping on validation MAE.
Stage 2: Evidential Training. The model is then fine-tuned using the evidential loss function $\mathcal{L}$, enabling uncertainty quantification while maintaining prediction accuracy. Learning rate reduced to $10^{-4}$, with gradient clipping (norm 1.0) to handle NIG sensitivity, epochs 20–50.
Stage 3: Calibration. Finally, the model undergoes calibration using temperature scaling [14] or isotonic regression on held-out data to ensure that uncertainty estimates are well-calibrated across different performance metrics and operating conditions (e.g., expected calibration error <0.05<0.05). This involves optimizing a post-hoc scalar per target to align predicted confidences with empirical accuracies.
Regularization. We apply graph edge dropout (p=0.1p=0.10.20.2), small feature noise ($\sigma=0.05$), multi-scale graph training, and mild device augmentation (e.g., $\pm 5\%$ DVFS tables). Scheduling-aware ranking losses further align predictions with downstream policy selection, yielding fast inference, cross-device transfer capability, and useful uncertainties for downstream schedulers and RL.
Ablation Summary. Table 3 quantifies the contribution of each architectural component to prediction accuracy. We systematically remove components from the full GraphPerf-RT model and measure the resulting degradation in $R^{2}$ on the held-out test set.
Table 3: Ablation study: contribution of each component to prediction accuracy. Full model achieves $R^{2}=0.97$.
Configuration $R^{2}$ $\Delta R^{2}$ MAPE
Full GraphPerf-RT 0.97 7.4%
     w/o Heterogeneous edges 0.86 -0.11 12.8%
     w/o Runtime telemetry 0.88 -0.09 11.5%
     w/o CFG features 0.91 -0.06 9.7%
     w/o Evidential heads (MSE only) 0.96 -0.01 7.9%
MLP baseline (no graph) 0.81 -0.16 15.2%
Homogeneous GCN 0.86 -0.11 12.4%
The heterogeneous edge types ($E_{TR}$, $E_{RR}$, $E_{RM}$) contribute the largest improvement (+11% $R^{2}$), confirming that modeling task-resource interactions explicitly outperforms homogeneous message passing. Runtime telemetry (thermal state, utilization, perf counters) adds +9% by capturing dynamic system behavior invisible to static analysis. CFG features contribute +6% by encoding control-flow complexity that predicts irregular execution patterns. Evidential heads minimally affect mean accuracy but are essential for calibrated uncertainty (see Section 5.3).
@@ 5.1 Experimental Setup
Hardware Platforms. We evaluate GraphPerf-RT on three embedded ARM SoCs that represent diverse heterogeneous architectures and thermal management capabilities.
Benchmark Suites. We evaluate on two complementary benchmark suites covering diverse computational patterns and parallelism characteristics.
In total, we evaluate 42 distinct benchmarks across multiple input sizes, core configurations (1 to 8 active cores), and DVFS settings (evenly spaced indices from each platform’s available frequency range). Both sequential and task-parallel execution modes are profiled, resulting in the 73,920 samples described in Section 3.3.
@@
Table 4: Overall performance across all platforms and benchmarks. Lower is better for RMSE/MAE/MAPE. Higher is better for $R^{2}$/Spearman.
Model RMSE MAE MAPE $R^{2}$ Spearman
Linear Regression 2.84 2.12 34.2% 0.67 0.71
Random Forest 2.31 1.78 28.5% 0.74 0.78
MLP 1.95 1.45 23.1% 0.81 0.83
GCN 1.67 1.23 19.8% 0.86 0.87
Het. Graph Trans. 1.52 1.11 17.3% 0.88 0.89
GraphPerf-RT 0.53 0.38 7.4% 0.97 0.95
Data Collection Infrastructure. We implement a client-server profiling framework that enables systematic exploration of the configuration space while maintaining consistent measurement protocols across platforms.
The host server orchestrates the experimental campaign by generating action tuples $a=(m,f,p)$ representing core mask, DVFS vector, and optional priority settings. Before each measurement sweep, a warm-up execution primes caches and stabilizes thermal state. Each completed run produces one CSV row containing timestamp, iteration index, execution mode, assigned and measured frequencies, active core configuration, input parameters, wall-clock execution time, per-rail energy and power readings, all performance counter values, and thermal zone temperatures. This unified logging schema enables direct comparison across TX2, Orin NX, and RUBIK Pi without format conversion, as detailed in Section 3.3.
Graph Extraction Pipeline. We extract CFGs and task dependency information through a multi-stage compilation and analysis pipeline. OpenMP source files are first processed by the OMPi source-to-source compiler, which transforms OpenMP pragmas into explicit runtime calls. The ALF-llvm backend then generates both LLVM intermediate representation (*.ll) and ARTIST2 Language for Flow analysis files (*.alf). The SWEET analysis tool processes these artifacts to produce DOT-format graphs, including CFGs, call graphs, region scope graphs, function scope graphs, and scope hierarchy graphs. A post-processing stage maps ALF entities to OpenMP tasks, merges chains of trivially small tasks while preserving dependency provenance, and computes topological encodings including depth, distance-to-sink, and centrality measures as described in Section 3.2.
5.2 Baselines and Metrics
Graph Convolutional Network (GCN) applies homogeneous message passing where all nodes and edges are treated uniformly, losing the type distinctions between task and resource nodes defined in Section 3.2.
Heterogeneous Graph Transformer (HGT) [16] represents a state-of-the-art heterogeneous GNN that uses type-aware attention over our node and edge schema as defined in Section 3.2. This baseline isolates the contribution of our evidential heads and runtime context integration, as HGT lacks evidential uncertainty quantification, thermal and utilization context in node features, and our CFG-derived task features.
5.3 Results and Analysis
Overall Prediction and Ranking Performance. Table 4 reports aggregate results across all platforms and benchmarks. GraphPerf-RT achieves the lowest prediction error and highest ranking quality among all evaluated methods. Compared to the strongest baseline (HGT), GraphPerf-RT reduces RMSE by 65% (from 1.52 to 0.53), MAE by 66% (from 1.11 to 0.38), and MAPE by 57% (from 17.3% to 7.4%). The $R^{2}$ value of 0.97 indicates that the model explains 97% of the variance in performance metrics, demonstrating strong predictive capability across diverse workloads and configurations.
@@ 5.3 Results and Analysis
The progression from tabular baselines (Linear Regression, Random Forest, MLP) to graph-based methods (GCN, HGT, GraphPerf-RT) reveals the importance of structural information. The MLP baseline, which uses identical features but ignores graph structure, achieves $R^{2}=0.81$, while the homogeneous GCN improves to $R^{2}=0.86$ by incorporating node connectivity. HGT further improves to $R^{2}=0.88$ through type-aware attention over heterogeneous edges. GraphPerf-RT achieves the best $R^{2}=0.97$ by combining heterogeneous message passing with evidential uncertainty heads, runtime context integration, and CFG-derived task features.
Platform-Specific Analysis. Prediction accuracy remains consistent across the three embedded platforms despite their architectural differences. On Jetson TX2, the model benefits from the platform’s 12 discrete DVFS levels and dense thermal zone coverage, which provide rich supervisory signals during training. Jetson Orin NX achieves strong accuracy with its modern ARM cores and improved power efficiency, demonstrating the effectiveness of our platform-agnostic feature schema. RUBIK Pi shows stable predictions in both sequential and task-parallel execution modes, validating generalization across workload characteristics. Per-device z-score normalization combined with device sheet broadcasting as described in Section 3 contributes to this cross-platform stability.
Uncertainty Calibration. The evidential regression heads provide well-calibrated uncertainty estimates across all target metrics, as shown in Table 5. ECE values below 0.05 for all metrics indicate that predicted confidence intervals contain the true value at approximately the stated rate. Makespan prediction achieves ECE of 0.043 with reliability of 0.956, meaning that 95.6% of predictions fall within their 95% confidence intervals. Energy prediction shows the best calibration with ECE of 0.038 and reliability of 0.962. Cache and branch miss predictions exhibit slightly higher ECE (0.051 and 0.047 respectively) due to the inherent variability in these microarchitectural metrics, but remain well-calibrated for practical use. Reliability curves closely track the identity line across all targets, confirming that the model’s confidence estimates are trustworthy for risk-aware scheduling decisions. Sharpness values indicate reasonably narrow predictive intervals, providing informative rather than overly conservative uncertainty bounds.
Table 5: Uncertainty calibration summary (lower ECE/MCE is better; higher reliability is better).
Metric ECE MCE Reliability Sharpness
Makespan 0.043 0.089 0.956 0.234
Energy 0.038 0.076 0.962 0.198
Cache Misses 0.051 0.102 0.949 0.267
Branch Misses 0.047 0.094 0.953 0.251
Ablation Studies. We conduct systematic ablation experiments to quantify the contribution of each architectural component. Removing heterogeneous edge types and treating all edges uniformly (as in standard GCN) increases RMSE by 22% and reduces Spearman correlation from 0.95 to 0.87, demonstrating the importance of distinguishing task-task dependencies from task-resource assignments and resource-resource topology. Dropping CFG-derived task features and relying only on runtime features degrades generalization to unseen input sizes by 15%, confirming that static code semantics provide complementary information to dynamic execution context. Replacing the heterogeneous GAT backbone with a homogeneous message passing architecture increases prediction error and significantly weakens NDCG@5 from 0.94 to 0.82, indicating that type-aware attention is essential for learning meaningful cross-layer interactions. Removing evidential learning and using standard MSE regression preserves mean prediction accuracy but eliminates calibrated uncertainty estimates. Without uncertainty gating, the proportion of unsafe scheduling proposals (those that would cause thermal violations) increases from 3% to 18%, demonstrating the practical value of calibrated confidence intervals for risk-aware decision making.
Computational Efficiency. GraphPerf-RT is designed for deployment on resource-constrained embedded systems where inference latency directly impacts scheduling responsiveness. On-device inference completes in 2 to 5 milliseconds for typical OpenMP task graphs, enabling real-time configuration evaluation during scheduling decisions. The computational cost per GAT layer is $O(H\cdot|E|\cdot d)$ where $H$ denotes attention heads, $|E|$ is the edge count, and $d$ is the hidden dimension. With our configuration of $L\leq 6$ layers, $d\leq 128$ dimensions, and typical graph sizes of hundreds of edges, total inference remains within millisecond budgets. Model memory footprint is approximately 15 to 25 MB depending on layer depth, fitting comfortably within the memory constraints of Jetson TX2, Orin NX, and RUBIK Pi platforms.
Cross-Platform Generalization. To evaluate transfer capabilities, we train GraphPerf-RT on data from two platforms and evaluate on the held-out third platform without fine-tuning. This leave-one-platform-out protocol tests whether learned representations capture fundamental software-hardware interactions that generalize beyond specific architectural details. Results show modest accuracy degradation (average RMSE increase of 18%) when transferring to unseen platforms, but predictions remain practically useful for scheduling guidance. Importantly, ranking quality degrades less than point prediction accuracy: Spearman correlation decreases by only 8% on average, preserving the model’s ability to correctly order candidate configurations. Device sheet broadcasting and per-device feature normalization mitigate domain shift by anchoring predictions to platform-specific DVFS ranges and thermal characteristics. These transfer results suggest that the heterogeneous graph representation captures portable performance relationships that extend beyond the training platforms.

### v3

@@ 4 Design Methodology
This section presents the GraphPerf-RT architecture. As shown in Fig. 3, the pipeline takes heterogeneous graph input with typed nodes ($V_{T}$ for tasks, $V_{R}$ for resources, $V_{M}$ for memory), processes it through a GNN encoder with typed message passing, and produces multi-task predictions with calibrated uncertainty via evidential regression heads. The model takes task DAG $G=(V,E)$, device sheet $D$, runtime state $\mathcal{S}$, and action $a=(m,f,p)$, outputting predictions $y$ with calibrated uncertainty $\mathcal{U}$.
Figure 3: GraphPerf-RT architecture: heterogeneous graph encoding to evidential multi-task prediction.
@@ 4.1 Heterogeneous Graph Representation
Appendix M.1
Node Types (Appendix M.2). Task nodes $V_{T}$ encode CFG-derived features (loop count, max depth, cyclomatic complexity, branch density), DAG topology metrics (depth, distance-to-sink, betweenness centrality), static code statistics (estimated instructions, bytes moved), and optional runtime snapshots (recent performance counters, run-mode flags). Resource nodes $V_{R}$ encode per-core state: DVFS step (index and one-hot), core mask bit, cluster ID for heterogeneous architectures, utilization (exponential moving average), thermal headroom ($T_{\max}-T_{\text{current}}$), temperature trend ($\Delta T$ over recent samples), and bandwidth proxy. Memory nodes $V_{M}$ encode cache hierarchy: level identifier, capacity/associativity/line size, and latency/bandwidth proxies.
@@
Device constants from $D$ (DVFS tables, cache sizes, governor flags) are broadcast as shared features during node encoding.
Edge Types (Appendix M.3). Four edge types capture performance-critical interactions. $E_{TT}$ (task-task) edges encode precedence constraints with attributes: dependency type (spawn/join/data), critical-edge flag (on longest path), hop distance, and contention proxies. The critical path determines minimum makespan; fork/join patterns reveal synchronization overhead. $E_{TR}$ (task-resource) edges connect tasks to cores under the scheduling assignment, with affinity strength and migration overhead attributes. Task-to-core affinity affects cache locality; load imbalance increases makespan. $E_{RR}$ (resource-resource) edges connect cores sharing hardware components (L2 caches, memory controllers) to model contention, with sharing degree and interconnect latency attributes. $E_{RM}$ (resource-memory) edges connect cores to cache levels, encoding bandwidth allocation. Memory-bound workloads are bottlenecked by bandwidth; cache miss rates determine effective memory latency.
4.2 GNN Architecture
Appendix M.4
@@ 4.2 GNN Architecture
Type-Specific Encoders. Since different node types have distinct feature spaces, we employ type-specific MLPs to encode raw features into a common embedding space of dimension $d=128$: $\mathbf{h}_{v}^{(0)}=\text{MLP}_{\text{type}(v)}(\mathbf{x}_{v};\theta_{\text{type}(v)})$ where $\text{type}(v)\in\{T,R,M\}$. Each MLP has 2–3 layers with ReLU activations and dropout $p=0.1$. Topological encodings (positional embeddings from DAG depth, distance-to-sink, core cluster membership) are concatenated before encoding.
Heterogeneous GAT Layers. The core architecture consists of 3–6 heterogeneous graph attention layers that aggregate information from neighbors while accounting for edge types. For each layer $l$: $\mathbf{h}_{v}^{(l+1)}=\sigma\left(\mathbf{h}_{v}^{(l)}+\sum_{r\in\mathcal{R}}\sum_{u\in\mathcal{N}_{r}(v)}\alpha_{uv,r}^{(l)}\mathbf{W}_{r}^{(l)}\mathbf{h}_{u}^{(l)}\right)$ where $\mathcal{R}=\{TT,TR,RR,RM\}$, $\mathcal{N}_{r}(v)$ is the neighborhood under edge type $r$, $\alpha_{uv,r}^{(l)}$ is the attention coefficient, $\mathbf{W}_{r}^{(l)}$ is a type-specific transform, and $\sigma$ is ELU. Attention coefficients incorporate edge features $\mathbf{e}_{uv,r}$ (e.g., contention proxies). Multi-head attention (4–8 heads) captures diverse interaction patterns.
Graph-Level Pooling. After the final layer $L$, hierarchical pooling aggregates per-type embeddings: $\mathbf{h}_{G}=\text{CONCAT}(\text{POOL}_{T},\text{POOL}_{R},\text{POOL}_{M})$ where pooling can be mean, max, or attention-based. This yields a fixed-size graph representation (256–512 dimensions) regardless of graph size, enabling batched inference.
@@ 4.3 Evidential Learning for Uncertainty Quantification
Instead of directly predicting performance values, GraphPerf-RT learns parameters of a Normal-Inverse-Gamma (NIG) distribution for each metric, enabling uncertainty-aware predictions in a single forward pass (Appendix M.5).
Prediction Heads. For each metric $k\in\{\text{makespan},\text{energy},\text{cache misses},\text{branch misses},\text{utilization}\}$, an evidential head outputs NIG parameters: $(\gamma_{k},\nu_{k},\alpha_{k},\beta_{k})=\text{MLP}_{k}(\mathbf{h}_{G};\theta_{k})$ with $\nu_{k}=\text{softplus}(\tilde{\nu}_{k})+1$, $\alpha_{k}=\text{softplus}(\tilde{\alpha}_{k})+1$, $\beta_{k}=\text{softplus}(\tilde{\beta}_{k})$ ensuring valid NIG parameters ($\nu_{k},\alpha_{k}>1,\beta_{k}>0$). The mean prediction is $\hat{y}_{k}=\gamma_{k}$, with uncertainty decomposition: $\text{Aleatoric}_{k}=\beta_{k}/(\alpha_{k}-1)$, $\text{Epistemic}_{k}=\beta_{k}/(\nu_{k}(\alpha_{k}-1))$. Aleatoric uncertainty captures irreducible noise from OS scheduling and co-runner interference; epistemic uncertainty reflects model uncertainty that decreases with more training data.
Loss Function. Training minimizes the NIG negative log marginal likelihood with non-saturating uncertainty regularization [36] to address evidence contraction on high-error samples, plus a ranking loss aligned to makespan-first scheduling objectives. The non-saturating term prevents the model from being overconfident on out-of-distribution inputs by penalizing high evidence when predictions are inaccurate, ensuring robust uncertainty estimates critical for thermal safety. The ranking loss operates on Lower Confidence Bounds (LCB = $\gamma-k\sigma$, where $\sigma=\sqrt{\beta(\nu+1)/(\nu(\alpha-1))}$ is the total predictive standard deviation) rather than point estimates, promoting risk-aware ordering that prioritizes configurations with both low predicted makespan and high confidence.
@@ 4.4 Training Procedure
GraphPerf-RT is trained using a multi-stage procedure designed for multi-task learning with uncertainty quantification (Appendix M.7).
Stage 1: Feature Learning. The model is first trained using a standard multi-task regression loss (MSE per target, weighted by inverse variance), focusing on prediction accuracy without uncertainty quantification. Optimizer: AdamW with learning rate $10^{-3}$, batch size 32–128 depending on graph size, epochs 50–100 with early stopping on validation MAE.
Stage 2: Evidential Training. The model is fine-tuned using the evidential loss function, enabling uncertainty quantification while maintaining prediction accuracy. Learning rate reduced to $10^{-4}$ with gradient clipping (norm 1.0) to handle NIG sensitivity, epochs 20–50.
Stage 3: Calibration. Post-hoc calibration on held-out data scales the predictive standard deviation $\sigma^{\prime}=\tau\cdot\sigma$ where $\tau$ is fit to achieve target PICP>>95% at 95% confidence. The scalar $\tau$ is optimized globally (across all platforms) to minimize $|\text{PICP}-0.95|$ on the validation set. The model intentionally produces wider intervals than strictly necessary, ensuring scheduling decisions never rely on overconfident predictions—essential for safety-critical embedded systems.
Regularization. Graph edge dropout (p=0.1p=0.10.20.2), small feature noise ($\sigma=0.05$), and mild device augmentation ($\pm 5\%$ DVFS tables) improve generalization. Scheduling-aware ranking losses align predictions with downstream policy selection.
4.5 Ablation Summary
Table 1: Ablation: component contributions to prediction accuracy (log-transformed makespan).
Configuration $R^{2}$ $\Delta R^{2}$ Spearman
Full GraphPerf-RT 0.81 0.95
w/o Hetero edges 0.77 -0.04 0.87
w/o Runtime telemetry 0.78 -0.03 0.89
w/o CFG features 0.79 -0.02 0.91
w/o Memory nodes ($V_{M}$) 0.79 -0.02 0.92
w/o Evidential (MSE) 0.80 -0.01 0.94
MLP (no graph) 0.75 -0.06 0.83
Homogeneous GCN 0.77 -0.04 0.87
Heterogeneous edges contribute the largest ranking improvement (Spearman +8%), followed by runtime telemetry (+6%), CFG features (+4%), and memory nodes (+3%). Memory nodes capture cache hierarchy interactions critical for memory-bound workloads. Evidential heads minimally affect point accuracy but are essential for calibrated uncertainty (PICP=99.9%), enabling safe scheduling decisions.
@@ 5.1 Experimental Setup
Hardware Platforms. We evaluate GraphPerf-RT on three embedded ARM SoCs representing diverse architectures (Appendix L.1). NVIDIA Jetson TX2 features a heterogeneous hexa-core configuration combining two Denver 2 cores with four Cortex-A57 cores, 12 discrete DVFS levels (345.6 MHz–2.0 GHz), and multiple thermal zones with a 50°C cap. RUBIK Pi is an 8-core Cortex-A72-class SBC with per-core userspace DVFS and energy measurement via qcom-battmgr or hwmon rails. NVIDIA Jetson Orin NX provides 8 Cortex-A78AE cores with userspace DVFS, representing the latest generation of embedded AI platforms. Across all platforms, we configure the Linux cpufreq subsystem for userspace governor control, discover DVFS indices dynamically, and sample thermal zones before/after execution.
Benchmark Suites. We evaluate 42 programs from two complementary suites (Appendix L.2). BOTS provides 12 task-parallel applications: alignment, fft, fib, floorplan, health, concom, knapsack, nqueens, sort, sparselu, strassen, and uts (unbalanced tree search with irregular parallelism). PolyBench contributes 30 kernels spanning linear algebra (gemm, cholesky, lu), stencils (jacobi-2d, seidel-2d), and data mining (correlation, covariance). Each benchmark executes with multiple inputs, core counts (1–8), and DVFS settings. Total dataset: 73,920 samples with 60/20/20 splits stratified by (benchmark, input, core mask, DVFS index).
Data Collection. We implement a client-server profiling framework (Appendix L.3). The device client configures cpufreq, applies DVFS indices, launches benchmarks via a real-time wrapper, and collects performance counters (cycles, instructions, cache/branch misses), power samples, and thermal readings. Each run produces one CSV row with timestamps, frequencies, execution time, energy, counters, and temperatures.
Graph Extraction. OpenMP sources are compiled through OMPi; ALF-llvm emits LLVM IR and ALF files; SWEET generates CFG/call graphs (Appendix L.4). Post-processing maps entities to OpenMP tasks with topological encodings (depth, distance-to-sink, centrality).
Baselines. We compare against seven approaches (Appendix L.5): Linear Regression (Ridge regularization on tabular features), Random Forest (100 trees), MLP (3-layer feedforward), GCN (homogeneous message passing), Heterogeneous Graph Transformer (HGT) with type-aware attention, plus two uncertainty quantification variants (HGT+Ensemble with 5 models, HGT+MC Dropout). All use identical splits and tuning protocols.
5.2 Results
@@
Table 2: Overall performance on log-transformed makespan. Lower RMSE/MAE better; higher $R^{2}$/Spearman better.
Model RMSE MAE $R^{2}$ Spearman
Linear Reg. 1.21 0.89 0.67 0.71
Random Forest 0.98 0.72 0.74 0.78
MLP 0.83 0.59 0.75 0.83
GCN 0.71 0.50 0.77 0.87
HGT 0.65 0.45 0.78 0.89
[…]
HGT + MC Dropout 0.64 0.44 0.78 0.89
GraphPerf-RT 0.45 0.24 0.81 0.95
[…]
@@
The progression from tabular baselines (Linear, RF, MLP) to graph-based methods (GCN, HGT, GraphPerf-RT) reveals the importance of structural information. MLP achieves Spearman=0.83 without graph structure; GCN improves to 0.87 with homogeneous edges; HGT reaches 0.89 with type-aware attention; GraphPerf-RT achieves 0.95 by combining heterogeneous message passing with evidential heads and runtime context.
Uncertainty Calibration. Table 3 shows uncertainty calibration metrics for makespan prediction. The model achieves exceptional Prediction Interval Coverage Probability (PICP) of 99.9% at 95% confidence, meaning the true values fall within the predicted intervals more often than the nominal rate. This conservative calibration is intentional and desirable for safety-critical embedded systems: the model produces slightly wider intervals than strictly necessary, ensuring scheduling decisions never rely on overconfident predictions. The uncertainty decomposition shows 94% aleatoric and 6% epistemic, indicating the model correctly identifies data-driven noise (OS scheduling jitter, thermal variation) as dominant, with model uncertainty contributing a smaller component that decreases with additional training data.
Table 3: Uncertainty calibration for makespan. MPIW = Mean Prediction Interval Width. PICP>>nominal indicates conservative calibration.
Confidence PICP MPIW Coverage Gap
95% 99.9% 58.1 +4.9%
90% 99.5% 47.2 +9.5%
80% 98.1% 35.8 +18.1%
Computational Efficiency. On-device inference completes in 2–7ms for typical task graphs (8 nodes, 56 edges), enabling real-time scheduling. Model size is 12.4MB, fitting embedded memory constraints (Appendix I.11).